OpenStudy (anonymous):

solve x^2-8x=-16 by completing the square. is it 4,4?

OpenStudy (anonymous):

(x-2)^2

Directrix (directrix):

Yes, you have a double root because the parabola is tangent to the x-axis at (4,0).

Directrix (directrix):

(x-2)^2 = ? @michaelvelarde

OpenStudy (anonymous):

well anything you do to the left you have to do to the right so x^2-4x+4=-16+4